OpenStudy (theopenstudyowl):

lim (2^x-1)/(x) x->0

OpenStudy (anonymous):

it is the derivative of \)2^x\) at \(x=0\)

OpenStudy (anonymous):

if you got that far

OpenStudy (theopenstudyowl):

\[\lim_{x \rightarrow 0} (2^x-1)/x\]

OpenStudy (anonymous):

L'hopital's rule

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

i mean sure you can, but this is the derivative of \(f(x)=2^x\) at \(x=0\) i.e. it is \(f'(0)\)

OpenStudy (anonymous):

you know what the derivative of \(y=2^x\) is?

OpenStudy (anonymous):

Hint: \[\frac{ d }{ dx } a^x = a^x lnx\]

OpenStudy (anonymous):

@iambatman hmmm